2 conductor insulated from each other by transferring electrons from one conductor to other. After 25×19^12 electrons have been transfered. The potential difference between the conductor is found to be 16V. The capacitance of the system is

Grade:12

1 Answers

Arun
6 years ago
Given that Number of electron = 25 × 19^12 Net charge Q = 25 × 19^12 × 1.6 × 10^–19  ∴ The net potential difference = 16 V. ∴ Capacitance – C = q/v Put the value of q and v and get your answer.  RegardsArun (askIITians forum expert)